# An LLL algorithm with symmetries

### v1

@@ 2.2. Set-up for the case $\mathbf{G}=\mathrm{SL}_{g}$
($2\leq j\leq g$) with $\mu_{i,j}=(e_{j},e_{i}^{\ast})_{H}/(e_{i}^{\ast},e_{i}^{\ast})_{H}$. We take $\mu_{i,j}=1$ and $\mu_{i,j}=0$ if $i>j$. We then have (cf. [Kna02, Ch. VI, §4, Example]):
@@
These two operations correspond, respectively, to acting on $X_{\mathrm{SL}_{g}}$ on the right by elements of the form $\gamma_{n}=u_{\alpha_{i,j}}(m)\in\mathbf{N}(\mathbb{Z})$, and by elements of the form $s_{i}$. It is easy to express the effect of the first operation in terms of the co-ordinates $(a,n)$: $(a,n)$ is replaced by $(a,n\gamma_{n})$. The essential content of this step is that the set $\omega\leq N$ of matrices whose above-diagonal entries have absolute value at most $1/2$ is a fundamental set for the right action of $\mathbf{N}(\mathbb{Z})$ on $N$. It is trickier to express the effect of swapping basis vectors in these co-ordinates. In [LLL82], this is computed by looking at the effect on the Gram–Schmidt process for the new basis; we will see below that it suffices instead to do a computation in the root $\mathrm{SL}_{2}$ corresponding to the chosen simple root.
@@ 2.3. What it means to be reduced
We now return to the setting of §2.1, so $\mathbf{G}$ is a split semisimple group with split maximal torus torus $\mathbf{T}$. We fix a set $\Delta$ of simple roots as in §2.1 and let $\mathbf{N}$ be the corresponding unipotent subgroup. We are nearly ready to say what it means for a point $H\in X_{G}$ to be reduced. This will depend on a choice of parameter $\delta\in(1/4,1)$ and also a fundamental set $\omega\subset N$ for the right action of $\mathbf{N}(\mathbb{Z})$.
@@ 2.5. The algorithm
We maintain the notation of §2.3, so $\mathbf{G}$ is a split semisimple group over $\mathbb{Z}$ with split maximal torus $\mathbf{T}$, $\Delta$ is a fixed set of simple roots, $\mathbf{N}$ is the corresponding maximal unipotent subgroup of $\mathbf{G}$, and $K\leq G$ is the maximal compact subgroup defined in §2.1. In addition, for each $\alpha\in\Delta$, we fix a basis vector $X_{\alpha}$ for ${\mathfrak{g}}_{\alpha}$ with corresponding root $\mathrm{SL}_{2}$ given by $\phi_{\alpha}:\mathrm{SL}_{2}\to\mathbf{G}$.
We now describe the main construction of this paper:
Algorithm 2.10.
Fix $\delta\in(1/4,1)$ and an allowable fundamental set $\omega\subset N$. Given elements $a\in A$, $n\in N$, this algorithm produces an element $\gamma\in\mathbf{G}(\mathbb{Z})$ such that $x=Kan\gamma\in X_{G}$ is $(\delta,\omega)$-reduced while also computing $a(x)\in A,n(x)\in N$ such that $x=Ka(x)n(x)$.
  1. (1)
    (Initialise) Set $\gamma\leftarrow 1\in\mathbf{G}(\mathbb{Z})$, $a^{\prime}\leftarrow a$, $n^{\prime}\leftarrow n$.
  2. (2)
    (Size-reduce) Find $\gamma_{n^{\prime}}\in\mathbf{N}(\mathbb{Z})$ such that $n^{\prime}\gamma_{n^{\prime}}\in\omega$. Set $\gamma\leftarrow\gamma\gamma_{n^{\prime}}$, $n^{\prime}\leftarrow n^{\prime}\gamma_{n^{\prime}}$.
  3. (3)
    (Test $\mathbf{G}$-Lovász condition) Find $\alpha\in\Delta$ such that $\alpha(a^{\prime})^{2}>(\delta-p_{\alpha}(n^{\prime})^{2})^{-1}$ (if one exists). If there is no such $\alpha$, go to Step 4. Otherwise, let $t=p_{\alpha}(n)$, $m=u_{\alpha}(-t)n$, and set
    $\gamma\leftarrow\gamma s_{\alpha},$
    $a^{\prime}\leftarrow\check{\alpha}\left(\sqrt{1+t^{2}\alpha(a^{\prime})^{2}}\right)s_{\alpha}(a^{\prime}),$
    and
    $n^{\prime}\leftarrow u_{\alpha}\left(\frac{-t\alpha(a^{\prime})^{2}}{1+t^{2}\alpha(a^{\prime})^{2}}\right)s_{\alpha}^{-1}ms_{\alpha},$
    then go to Step 2.
  4. (4)
    (End) Return $\gamma$, $a^{\prime}$, and $n^{\prime}$. Then if $x=Kan\gamma$, we may take $a(x)=a^{\prime}$ and $n(x)=n^{\prime}$.
@@
Let $\kappa$ denote the Killing form of ${\mathfrak{g}}_{\mathbb{R}}$, and let $(\cdot,\cdot)_{0}$ denote the inner product $(X,Y)_{0}=-\kappa(X,d\theta_{0}Y)$ on ${\mathfrak{g}}_{\mathbb{R}}$ (it is positive definite, and therefore an inner product, because $\theta_{0,\mathbb{R}}$ is a Cartan involution.) If $g\in G$, then we write $(\cdot,\cdot)_{g}$ for the inner product $(X,Y)_{g}=(\operatorname{Ad}(g)(X),\operatorname{Ad}(g)(Y))_{0}$. Thus $(\cdot,\cdot)_{g}$ only depends on the image of $g$ in $X_{G}=K\backslash G$. We observe that $(\cdot,\cdot)_{0}$ and $(\cdot,\cdot)_{g}$ determine inner products on the exterior powers $\wedge^{r}{\mathfrak{g}}_{\mathbb{R}}$ in a natural way. We will also (by abuse of notation) denote these by $(\cdot,\cdot)_{0}$ and $(\cdot,\cdot)_{g}$, respectively.
Let ${\mathfrak{b}}=\mathfrak{t}\oplus\sum_{\alpha\in\Phi^{+}}\mathfrak{g}_{\alpha}$, and let $\ell=\dim{\mathfrak{b}}$. Then $\wedge^{\ell}{\mathfrak{b}}$ is a rank-one $\mathbb{Z}$-submodule of $\wedge^{\ell}\mathfrak{g}$. Choose $b\in\wedge^{\ell}{\mathfrak{b}}$ to be a basis vector. We can compute for any $a_{1}\in A$, $n_{1}\in N$:
@@
We now observe that $\operatorname{Ad}(\gamma_{1})b$ is another non-zero vector in the lattice $\wedge^{\ell}{\mathfrak{g}}$. Let $c_{1}=(b,b)_{an}^{1/2}$ denote the length of $b$ with respect to the inner product $(\cdot,\cdot)_{an}$, and let $c_{0}$ denote the minimal length of a non-zero vector in $\wedge^{\ell}{\mathfrak{g}}$, with respect to the inner product $(\cdot,\cdot)_{an}$. We finally obtain the inequality $c_{0}<\delta^{k}c_{1}$, hence $k<\log(c_{0}/c_{1})/\log(\delta)=\log(c_{1}/c_{0})/\log(\delta^{-1})$. The right-hand side here only depends on $an$ and $\delta$, showing that the algorithm must terminate after at most $\lfloor\log(c_{1}/c_{0})/\log(\delta^{-1})\rfloor$ iterations of Step 3. ∎
@@ 3. The case $\mathbf{G}=\operatorname{Sp}_{2g}$
If $-g\leq i\leq g$ and $i\neq 0$, let $\varepsilon_{i}\in X^{\ast}(\mathbf{T})$ be the character sending $\operatorname{diag}(t_{-g},\dots,t_{g})$ to $t_{i}$. The roots $\Phi=\Phi(\operatorname{Sp}_{2g},\mathbf{T})$ are given by $\varepsilon_{i}-\varepsilon_{j}$ ($-g\leq i,j\leq-1$, $i\neq j$) and ±$\pm(\varepsilon_{i}+\varepsilon_{j})$ ($-g\leq i\leq j\leq-1$). A system of positive roots is $\Phi^{+}=\{\varepsilon_{i}-\varepsilon_{j}\mid-g\leq i<j\leq-1\}\cup\{\varepsilon_{i}+\varepsilon_{j}\mid-g\leq i\leq j\leq-1\}$. The corresponding Borel subgroup $\mathbf{B}\subset\operatorname{Sp}_{2g}$ is the subgroup of upper-triangular matrices preserving $S$; its unipotent radical $\mathbf{N}$ is the subgroup of unipotent upper-triangular matrices preserving $S$. The group $A=\mathbf{T}(\mathbb{R})^{\circ}$ is the subgroup of diagonal matrices with positive entries preserving $S$.
@@
We first describe the (right) actions of $\operatorname{Sp}_{2g}(\mathbb{R})$ on the sets (1) – (3). The action on $X_{\operatorname{Sp}_{2g}(\mathbb{R})}$ is the usual one on right cosets. If $x\in\operatorname{Sp}_{2g}(\mathbb{R})$ and $H$ is an inner product as in (2), then we define $H\cdot x={}^{t}xHx$. If $J$ is in (3), we define $J\cdot x=x^{-1}Jx$.
Let $H_{0}$ denote the standard inner product on $\mathbb{R}^{2g}$. Given $x\in\operatorname{Sp}_{2g}(\mathbb{R})$, we associate the inner product ${}^{t}xH_{0}x$. The stabiliser equals $K$, so this gives an injective, $\operatorname{Sp}_{2g}(\mathbb{R})$-equivariant map $(1)\rightarrow(2)$. Thus the map is bijective. Given $H$ in $(2)$, let $J=-S^{-1}H$. An elementary computation shows that $J$ lies in (3), and that the map $H\mapsto J$ is $\operatorname{Sp}_{2g}(\mathbb{R})$-equivariant. An inverse is given by $H=-SJ$. Thus (2) and (3) are in bijection.
@@
with $\mu_{i,j}=(e_{j},e_{i}^{\ast})_{H}/(e_{i}^{\ast},e_{i}^{\ast})_{H}$. In particular, we take $\mu_{i,j}=1$ and $\mu_{i,j}=0$ if $i>j$.
@@
  • For each $i=-g+1,\dots,-1$, successively make the entries $n_{j,i}$ ($j=i-1,\dots,-g$) lie in $[-1/2,1/2)$ by multiplying on the right by $u_{\varepsilon_{i}-\varepsilon_{j}}(m_{i,j})$ for some integer $m_{i,j}$.
  • For each $i=-1,\dots,-g$, successively make the entries $n_{j,-i}$ ($j=i,\dots,-g$) lie in $[-1/2,1/2)$ by multiplying on the right by $u_{\varepsilon_{i}+\varepsilon_{j}}(m_{i,j})$ for some integer $m_{i,j}$.
@@
We next make explicit the simple reflections $s_{-g},\dots,s_{-1}$ respectively corresponding to the chosen simple roots $\varepsilon_{i}-\varepsilon_{i+1}$ ($i=-g,\dots,-2$), $2e_{-1}$, by describing their action on basis vectors:
@@
Algorithm 3.3.
@@
Fix $\delta\in(1/4,1)$. This algorithm takes as input an inner product $H$ on $\mathbb{R}^{2g}$ compatible with $S$, and produces an element $\gamma\in\operatorname{Sp}_{2g}(\mathbb{Z})$ such that ${}^{t}\gamma H\gamma$ corresponds to a $(\delta,\omega)$-reduced point of $X_{\operatorname{Sp}_{2g}(\mathbb{R})}$ under the bijection of Proposition 3.1.
  1. (1)
@@
    (Initialise) Using Gram–Schmidt, compute $a\in A,n\in N$ such that $H={}^{t}(an)an$. Set $a^{\prime}\leftarrow a$, $n^{\prime}\leftarrow n$, $\gamma\leftarrow 1\in\operatorname{Sp}_{2g}(\mathbb{Z})$, $k\leftarrow-g$.
  2. (2)
    (Iteration for $k<-1$) Execute RED($k,k+1$).
    1. (a)
      If $\alpha_{k}(a^{\prime})^{2}>(\delta-(n^{\prime}_{k,k+1})^{2})^{-1}$: execute REFL($k$). If $k>-g$, set $k\leftarrow k-1$ and go to Step 2. Otherwise, go to Step 2.
    2. (b)
      If $\alpha_{k}(a^{\prime})^{2}\leq(\delta-(n^{\prime}_{k,k+1})^{2})^{-1}$: execute RED($i,k+1$) for $i=k-1,\dots,-g$. Set $k\leftarrow k+1$. If $k<-1$, go to Step 2. If $k=-1$, go to Step 3.
  3. (3)
    (Iteration for $k=-1$) Execute RED($-1,1$).
    1. (a)
      If $\alpha_{k}(a^{\prime})^{2}>(\delta-(n^{\prime}_{-1,1})^{2})^{-1}$: execute REFL($-1$). Set $k\leftarrow k-1$ and go to Step 2.
    2. (b)
      If $\alpha_{k}(a^{\prime})^{2}\leq(\delta-(n^{\prime}_{-1,1})^{2})^{-1}$: execute RED($i,1$) for each $i=-2,\dots,-g$. For each $i=-2,\dots,-g$, execute RED($i,j$) for each $j=-i,\dots,-g$. Go to Step 4.
  4. (4)
    (End) Return $\gamma$.
  5. RED($i,j$)
    1. (a)
@@ 5. RED($i,j$)
      If $-g\leq i<j\leq-1$: let $m$ be the unique integer such that $n^{\prime}_{i,j}+t\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow nu_{\varepsilon_{i}-\varepsilon_{j}}(m)$, $\gamma\leftarrow\gamma u_{\varepsilon_{i}-\varepsilon_{j}}(m)$.
    2. (b)
@@
      If $-g\leq i\leq-1$ and $j=-i$: let $m$ be the unique integer such that $n^{\prime}_{i,j}+t\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow nu_{2\varepsilon_{i}}(m)$, $\gamma\leftarrow\gamma u_{2\varepsilon_{i}}(m)$.
    3. (c)
      If $-g\leq i<-j\leq-1$: let $m$ be the unique integer such that $n^{\prime}_{i,j}+t\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow nu_{\varepsilon_{i}+\varepsilon_{-j}}(m)$, $\gamma\leftarrow\gamma u_{\varepsilon_{i}+\varepsilon_{-j}}(m)$.
  6. REFL($i$)
    1. (a)
      If $-g\leq i\leq-2$: set $t\leftarrow n_{i,i+1}$, $m^{\prime}\leftarrow u_{\varepsilon_{i}-\varepsilon_{i+1}}(t)^{-1}n^{\prime}$, $n^{\prime}\leftarrow u_{\varepsilon_{i}-\varepsilon_{i+1}}((-t\alpha_{i}(a^{\prime})^{2})/(1+t^{2}\alpha_{i}(a^{\prime})^{2}))s_{i}^{-1}m^{\prime}s_{i}$, $a^{\prime}\leftarrow\check{\alpha}_{i}(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}})w_{i}(a^{\prime})$, $\gamma\leftarrow\gamma s_{i}$.
    2. (b)
      If $i=-1$: set $t\leftarrow n_{-1,1}$, $m^{\prime}\leftarrow u_{2\varepsilon_{i}}(t)^{-1}n^{\prime}$, $n^{\prime}\leftarrow u_{2\varepsilon_{i}}((-t\alpha_{i}(a^{\prime})^{2})/(1+t^{2}\alpha_{i}(a^{\prime})^{2}))s_{i}^{-1}m^{\prime}s_{i}$, $a^{\prime}\leftarrow\check{\alpha}_{i}(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}})w_{i}(a^{\prime})$, $\gamma\leftarrow\gamma s_{i}$.

### v2

@@ 2.2. Set-up for the case $\mathbf{G}=\mathrm{SL}_{g}$
($2\leq j\leq g$) with $\mu_{i,j}=(e_{j},e_{i}^{\ast})_{H}/(e_{i}^{\ast},e_{i}^{\ast})_{H}$. We take $\mu_{i,i}=1$ and $\mu_{i,j}=0$ if $i>j$. We then have (cf. [Kna02, Ch. VI, §4, Example]):
@@
These two operations correspond, respectively, to acting on $X_{\mathrm{SL}_{g}}$ on the right by elements of the form $\gamma=u_{\alpha_{i,j}}(m)\in\mathbf{N}(\mathbb{Z})$, and by elements of the form $s_{i}$. It is easy to express the effect of the first operation in terms of the co-ordinates $(a,n)$: $(a,n)$ is replaced by $(a,n\gamma)$. The essential content of this step is that the set $\omega\leq N$ of matrices whose above-diagonal entries have absolute value at most $1/2$ is a fundamental set for the right action of $\mathbf{N}(\mathbb{Z})$ on $N$. It is trickier to express the effect of swapping basis vectors in these co-ordinates. In [LLL82], this is computed by looking at the effect on the Gram–Schmidt process for the new basis; we will see below that it suffices instead to do a computation in the root $\mathrm{SL}_{2}$ corresponding to the chosen simple root.
@@ 2.3. What it means to be reduced
We now return to the setting of §2.1, so $\mathbf{G}$ is a split semisimple group with split maximal torus $\mathbf{T}$. We fix a set $\Delta$ of simple roots as in §2.1 and let $\mathbf{N}$ be the corresponding unipotent subgroup. We are nearly ready to say what it means for a point $H\in X_{G}$ to be reduced. This will depend on a choice of parameter $\delta\in(1/4,1)$ and also a fundamental set $\omega\subset N$ for the right action of $\mathbf{N}(\mathbb{Z})$.
@@ 2.5. The algorithm
We maintain the notation of §2.3, so $\mathbf{G}$ is a split semisimple group over $\mathbb{Z}$ with split maximal torus $\mathbf{T}$, $\Delta$ is a fixed set of simple roots, $\mathbf{N}$ is the corresponding maximal unipotent subgroup of $\mathbf{G}$, and $K\leq G$ is the maximal compact subgroup defined in §2.1. In addition, we fix a labelling $\Delta=\{\alpha_{1},\dots,\alpha_{r}\}$ for the simple roots and, for each $\alpha\in\Delta$, we fix a basis vector $X_{\alpha}$ for ${\mathfrak{g}}_{\alpha}$ with corresponding root $\mathrm{SL}_{2}$ given by $\phi_{\alpha}:\mathrm{SL}_{2}\to\mathbf{G}$. Algorithm 2.10 describes the main construction of this paper.
Algorithm 2.10 Lattice reduction for $\mathbf{G}$.
Input: A constant $\delta\in(1/4,1)$, an allowable fundamental set $\omega\subset N$, and elements $a\in A$, $n\in N$. Output: An element $\gamma\in\mathbf{G}(\mathbb{Z})$ such that $x=Kan\gamma\in X_{G}$ is $(\delta,\omega)$-reduced, and elements $a(x)\in A$, $n(x)\in N$ such that $x=Ka(x)n(x)$.
1:  Set $\gamma\leftarrow 1\in\mathbf{G}(\mathbb{Z})$, $a^{\prime}\leftarrow a$, $n^{\prime}\leftarrow n$.
2:   Find $\gamma_{n^{\prime}}\in\mathbf{N}(\mathbb{Z})$ such that $n^{\prime}\gamma_{n^{\prime}}\in\omega$. Set $\gamma\leftarrow\gamma\gamma_{n^{\prime}}$, $n^{\prime}\leftarrow n^{\prime}\gamma_{n^{\prime}}$.
3:  for $i=1$ to $r$ do
4:     if $\alpha_{i}(a^{\prime})^{2}>(\delta-p_{\alpha_{i}}(n^{\prime})^{2})^{-1}$ then
5:        Set $t\leftarrow p_{\alpha_{i}}(n)$, $m\leftarrow u_{\alpha_{i}}(-t)n$, $\gamma\leftarrow\gamma s_{\alpha_{i}},$ $a^{\prime}\leftarrow\check{\alpha_{i}}\left(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}}\right)s_{\alpha_{i}}(a^{\prime}),$ and $n^{\prime}\leftarrow u_{\alpha_{i}}\left(\frac{-t\alpha_{i}(a^{\prime})^{2}}{1+t^{2}\alpha_{i}(a^{\prime})^{2}}\right)s_{\alpha_{i}}^{-1}ms_{\alpha_{i}}$.
6:        go to 2.
7:     end if
8:  end for
9:  Set $a(x)\leftarrow a^{\prime}$, $n(x)\leftarrow n^{\prime}$.
10:  return  $\gamma,a(x),n(x)$.
@@
Let $\kappa$ denote the Killing form of ${\mathfrak{g}}_{\mathbb{R}}$, and let $(\cdot,\cdot)_{0}$ denote the inner product $(X,Y)_{0}=-\kappa(X,d\theta_{0}Y)$ on ${\mathfrak{g}}_{\mathbb{R}}$ (it is positive definite, and therefore an inner product, because $\theta_{0,\mathbb{R}}$ is a Cartan involution.) If $g\in G$, then we write $(\cdot,\cdot)_{g}$ for the inner product $(X,Y)_{g}=(\operatorname{Ad}(g)(X),\operatorname{Ad}(g)(Y))_{0}$. Thus $(\cdot,\cdot)_{g}$ only depends on the image of $g$ in $X_{G}=K\backslash G$. We observe that $(\cdot,\cdot)_{0}$ and $(\cdot,\cdot)_{g}$ determine inner products on the exterior powers $\bigwedge^{r}{\mathfrak{g}}_{\mathbb{R}}$ in a natural way. We will also (by abuse of notation) denote these by $(\cdot,\cdot)_{0}$ and $(\cdot,\cdot)_{g}$, respectively.
Let ${\mathfrak{b}}=\mathfrak{t}\oplus\sum_{\alpha\in\Phi^{+}}\mathfrak{g}_{\alpha}$, and let $\ell=\dim{\mathfrak{b}}$. Then $\bigwedge^{\ell}{\mathfrak{b}}$ is a rank-one $\mathbb{Z}$-submodule of $\bigwedge^{\ell}\mathfrak{g}$. Choose $b\in\bigwedge^{\ell}{\mathfrak{b}}$ to be a basis vector. We can compute for any $a_{1}\in A$, $n_{1}\in N$:
@@
We now observe that $\operatorname{Ad}(\gamma_{1})b$ is another non-zero vector in the lattice $\bigwedge^{\ell}{\mathfrak{g}}$. Let $c_{1}=(b,b)_{an}^{1/2}$ denote the length of $b$ with respect to the inner product $(\cdot,\cdot)_{an}$, and let $c_{0}$ denote the minimal length of a non-zero vector in $\bigwedge^{\ell}{\mathfrak{g}}$, with respect to the inner product $(\cdot,\cdot)_{an}$. We finally obtain the inequality $c_{0}<\delta^{k}c_{1}$, hence $k<\log(c_{0}/c_{1})/\log(\delta)=\log(c_{1}/c_{0})/\log(\delta^{-1})$. The right-hand side here only depends on $an$ and $\delta$, showing that the algorithm must terminate after at most $\lfloor\log(c_{1}/c_{0})/\log(\delta^{-1})\rfloor$ iterations of Step 3. ∎
@@ 3. The case $\mathbf{G}=\operatorname{Sp}_{2g}$
If $-g\leq i\leq g$ and $i\neq 0$, let $\varepsilon_{i}\in X^{\ast}(\mathbf{T})$ be the character sending $\operatorname{diag}(t_{-g},\dots,t_{g})$ to $t_{i}$. The roots $\Phi=\Phi(\operatorname{Sp}_{2g},\mathbf{T})$ are given by $\varepsilon_{i}-\varepsilon_{j}$ ($-g\leq i,j\leq-1$, $i\neq j$) and ±$\pm(\varepsilon_{i}+\varepsilon_{j})$ ($-g\leq i\leq j\leq-1$). A system of positive roots is $\Phi^{+}=\{\varepsilon_{i}-\varepsilon_{j}\mid-g\leq i<j\leq-1\}\cup\{\varepsilon_{i}+\varepsilon_{j}\mid-g\leq i\leq j\leq-1\}$. The corresponding set of simple roots is $\Delta=\{\varepsilon_{i}-\varepsilon_{i+1}\mid-g\leq i\leq-2\}\cup\{2\varepsilon_{-1}\}$. The corresponding Borel subgroup $\mathbf{B}\subset\operatorname{Sp}_{2g}$ is the subgroup of upper-triangular matrices preserving $S$; its unipotent radical $\mathbf{N}$ is the subgroup of unipotent upper-triangular matrices preserving $S$. The group $A=\mathbf{T}(\mathbb{R})^{\circ}$ is the subgroup of diagonal matrices with positive entries preserving $S$.
@@
We first describe the (right) actions of $\operatorname{Sp}_{2g}(\mathbb{R})$ on the sets (1) – (3). The action on $X_{\operatorname{Sp}_{2g}(\mathbb{R})}$ is the usual one on right cosets. If $x\in\operatorname{Sp}_{2g}(\mathbb{R})$ and $H$ is an inner product as in (2), then we define $H\cdot x={}^{t}xHx$. If $J$ is as in (3), we define $J\cdot x=x^{-1}Jx$.
Let $H_{0}$ denote the standard inner product on $\mathbb{R}^{2g}$. Given $x\in\operatorname{Sp}_{2g}(\mathbb{R})$, we associate the inner product ${}^{t}xH_{0}x$. The stabiliser equals $K$, so this gives an injective, $\operatorname{Sp}_{2g}(\mathbb{R})$-equivariant map $(1)\rightarrow(2)$. The argument of [Gra84, Lemma 7.11] shows that $\operatorname{Sp}_{2g}(\mathbb{R})$ acts transitively on the set of inner products in (2). Thus the map is bijective. Given $H$ in $(2)$, let $J=-S^{-1}H$. An elementary computation shows that $J$ lies in (3), and that the map $H\mapsto J$ is $\operatorname{Sp}_{2g}(\mathbb{R})$-equivariant. An inverse is given by $H=-SJ$. Thus (2) and (3) are in bijection.
@@
with $\mu_{i,j}=(e_{j},e_{i}^{\ast})_{H}/(e_{i}^{\ast},e_{i}^{\ast})_{H}$. In particular, we take $\mu_{i,i}=1$ and $\mu_{i,j}=0$ if $i>j$.
@@
  • For each $j=-g+1,-g+2,\dots,-2,-1$, successively make the entries $n_{i,j}$ (for $i=j-1,j-2,\dots,-g+1,-g$) lie in $[-1/2,1/2)$ by multiplying on the right by $u_{\varepsilon_{i}-\varepsilon_{j}}(m_{i,j})$ for some integer $m_{i,j}$.
  • For each $j=1,2,\dots,g$, successively make the entries $n_{i,j}$ (for $i=-j,-j-1,\dots,-g$) lie in $[-1/2,1/2)$ by multiplying on the right by $u_{\varepsilon_{i}+\varepsilon_{-j}}(m_{i,j})$ for some integer $m_{i,j}$.
@@
We next make explicit the simple reflections $s_{-g},\dots,s_{-1}$ respectively corresponding to the chosen simple roots $\varepsilon_{i}-\varepsilon_{i+1}$ ($i=-g,\dots,-2$), $2\varepsilon_{-1}$, by describing their action on basis vectors:
@@
Algorithm 3.3 Lattice reduction for $\operatorname{Sp}_{2g}$.
Input: A constant $\delta\in(1/4,1)$, and a positive definite $2g\times 2g$ real symmetric matrix $H$ compatible with $S$.
@@
Output: An element $\gamma\in\operatorname{Sp}_{2g}(\mathbb{Z})$ such that ${}^{t}\gamma H\gamma$ corresponds to a $(\delta,\omega)$-reduced point of $X_{\operatorname{Sp}_{2g}(\mathbb{R})}$ under the bijection of Proposition 3.1.
@@
1:  Using Gram–Schmidt, compute $a\in A,n\in N$ such that $H={}^{t}(an)an$. Set $\gamma\leftarrow 1\in\operatorname{Sp}_{2g}(\mathbb{Z})$, $a^{\prime}\leftarrow a$, $n^{\prime}\leftarrow n$, $k\leftarrow-g$.
2:  while $k<-1$ do
3:     Execute RED($k,k+1$).
4:     if  $\alpha_{k}(a^{\prime})^{2}>(\delta-(n^{\prime}_{k,k+1})^{2})^{-1}$  then
5:        Execute REFL($k$).
6:        if $k>-g$ then
7:           $k\leftarrow k-1$.
8:        end if
9:     else
10:        for  $i=k-1$, $i\geq-g$, $i\leftarrow i-1$  do
11:           Execute RED($i,k+1$).
12:        end for
13:        Set $k\leftarrow k+1$.
14:     end if
15:  end while
16:  Execute RED($-1,1$).
17:  if $\alpha_{-1}(a^{\prime})^{2}>(\delta-(n^{\prime}_{-1,1})^{2})^{-1}$ then
18:     Execute REFL($-1$). Set $k\leftarrow k-1$.
19:     go to 2.
20:  else
21:     for $i=-2$, $i\geq-g$, $i\leftarrow i-1$ do
22:        Execute RED($i,1$).
23:     end for
24:     for $j=2$, $j\leq g$, $j\leftarrow j+1$ do
25:        for $i=-j$, $i\geq-g$, $i\leftarrow i-1$ do
26:           Execute RED($i,j$).
27:        end for
28:     end for
29:  end if
30:  return  $\gamma$.
Algorithm 3.4 Subroutine RED($i,j$) for Algorithm 3.3.
  if $-g\leq i<j\leq-1$ then
@@
     Let $m$ be the unique integer such that $n^{\prime}_{i,j}+m\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow nu_{\varepsilon_{i}-\varepsilon_{j}}(m)$, $\gamma\leftarrow\gamma u_{\varepsilon_{i}-\varepsilon_{j}}(m)$.
  else if $-g\leq i\leq-j\leq-1$ then
@@
     Let $m$ be the unique integer such that $n^{\prime}_{i,j}+m\in[-1/2,1/2)$. Set $n^{\prime}\leftarrow nu_{\varepsilon_{i}+\varepsilon_{-j}}(m)$, $\gamma\leftarrow\gamma u_{\varepsilon_{i}+\varepsilon_{-j}}(m)$.
  end if
Algorithm 3.5 Subroutine REFL($i$) for Algorithm 3.3.
  if $-g\leq i\leq-2$ then
     Set $t\leftarrow n_{i,i+1}$, $m^{\prime}\leftarrow u_{\varepsilon_{i}-\varepsilon_{i+1}}(t)^{-1}n^{\prime}$, $n^{\prime}\leftarrow u_{\varepsilon_{i}-\varepsilon_{i+1}}((-t\alpha_{i}(a^{\prime})^{2})/(1+t^{2}\alpha_{i}(a^{\prime})^{2}))s_{i}^{-1}m^{\prime}s_{i}$, $a^{\prime}\leftarrow\check{\alpha}_{i}(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}})s_{i}(a^{\prime})$, $\gamma\leftarrow\gamma s_{i}$.
  else if $i=-1$ then
     Set $t\leftarrow n_{-1,1}$, $m^{\prime}\leftarrow u_{2\varepsilon_{i}}(t)^{-1}n^{\prime}$, $n^{\prime}\leftarrow u_{2\varepsilon_{i}}((-t\alpha_{i}(a^{\prime})^{2})/(1+t^{2}\alpha_{i}(a^{\prime})^{2}))s_{i}^{-1}m^{\prime}s_{i}$, $a^{\prime}\leftarrow\check{\alpha}_{i}(\sqrt{1+t^{2}\alpha_{i}(a^{\prime})^{2}})s_{i}(a^{\prime})$, $\gamma\leftarrow\gamma s_{i}$.
  end if